OpenStudy (anonymous):

Solve the following system. y = x + 3 2x + y = 9 (2, 5) (5, 2) (−2, 5) (2, −5)

OpenStudy (anonymous):

substitute x+3 from the first equation in for y in the second equation and solve. 2x+(x+3)=9 3x+3=9 3x=6 x=2 Use this value of x to solve for y in the first equation. y=2+3 y=5 Answer is (2,5)

OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 3 hours with a 30 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. d = 3(r − 30) and d = 2(r + 30) d = 3(r + 30) and d = 2(r − 30) d = 5(r + 30) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

[6.06] Which point lies in the solution set for the following system of inequalities? 2x + y < 4 x − y > 3 (2, −3) (3, 4) (5, −2) (0, 0)

OpenStudy (anonymous):

The graph of the following system of equations is −2x + y = 3 −4x + 2y = 6 Overlapping lines Parallel lines Intersecting lines

OpenStudy (anonymous):

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 11 3x + 4y = 24